# The Minimal Attached Eddy in Wall Turbulence: Statistical Foundations, Inverse Identification and Influence Kernels

### v4

@@ Abstract
Townsend’s attached eddy hypothesis models the logarithmic region of high Reynolds number wall turbulence as a random superposition of wall-attached, geometrically self-similar eddies whose sizes obey a scale-invariant population law. Building on the statistical framework of woodcock2015statistical, the present work (i) poses an inverse problem to infer the ideal single-eddy contribution (influence) functions for the mean velocity and Reynolds stresses from DNS moments, (ii) uses these inferred kernels to guide a minimal Biot–Savart-consistent hairpin-type eddy built from Rankine vortex rods together with an inviscid image system, and (iii) introduces a spectral Influence kernel that maps a self-similar eddy footprint to its one-dimensional energy spectrum. The Influence-kernel viewpoint yields a transparent explanation for the emergence (and limitations) of the linear part of the energy spectrum, provides a clear scale-by-scale decomposition and helps rationalize why simple eddy templates can reproduce a broad set of log-layer statistics once the mean-flow anchoring is fixed.
@@
Characterizing and predicting the turbulent flow in the vicinity of a wall continues to not only be a problem of scientific interest, but also one of great practical relevance. Significant knowledge and insight has been gained from the development of experimental diagnostics in the 1960s to the emergence of direct numerical simulations in the 1980s, sophisticated measurements in the 1990s, an emphasis on coherent structures in the 2000s and operator-based dynamical models over the past decade. Against this backdrop of increasingly sophisticated experimentation, computations and analyses, the attached eddy hypothesis - originally proposed more than 50 years ago by  townsend1976structure - continues to serve as a simple, yet effective theory to qualitatively and quantitatively describe structural and statistical aspects of turbulent boundary layers, particularly focusing on the logarithmic region. The attached eddy model (AEM) was given a firm mathematical footing by  perry1982mechanism, and further developed by Marusic and co-workers over the past 25 years  (perry1995wall; marusic2013logarithmic; woodcock2015statistical; de2016influence). Among other useful features, the AEM is able to explain scaling behaviors of velocity moments, provide an explanation for uniform momentum zones (de2016uniform), and serve as a predictive model for the von-Karman constant as a function of Reynolds number. A review of the theory and developments of the AEM of wall turbulence can be found in  marusic2019attached. Over the past few years, theoretical and numerical analyses (e.g.  mckeon2019self; lozano2019characteristic) have added further credibility to this theory. Despite the success of the AEM, it is pertinent to remember that it is fundamentally a statistical theory, and other hypotheses (e.g  davidson2006logarithmic; davidson2009simple) can also be used to explain the statistical characteristics of the log layer.
@@
A key advance was made by woodcock2015statistical (henceforth W&M) who established a statistical foundation for AEM. They provide a complete derivation for all the velocity moments and demonstrate logarithmic scaling relationships therein. They were also able to provide expressions for the skewness and flatness of the wall-normal and spanwise fluctuations as a function of the Reynolds number. While variants of the AEM continue to be developed in the literature (e.g. hwang2020attached), we consider W&M as the starting point of our exploration and exclusively consider zero pressure gradient boundary layers. The main contributions and guiding questions in this work are:
@@
    Inverse identification of attached-eddy influence functions. Townsend introduced the notion of an “eddy intensity” (or contribution) function: a wall-parallel average of the single-eddy induced velocity and its products. While Townsend gave remarkably insightful descriptions of the nature of this function (Figure 5.7, page 155 of townsend1976structure), his and most other treatments are qualitative. Here we pose a concrete inverse problem that infers the ideal influence functions $I_{1}(y/h)$ and $I_{ij}(y/h)$ implied by reference one-point moments (DNS/experiment). The inferred kernels expose the wall-normal support, sign structure, and plateau/decay behavior required for log-layer mean and Reynolds stresses, and motivate a simplified kernel model that yields a direct relation between the von Kármán constant $\kappa$ and the population density $\beta$.
@@
    A minimal Biot–Savart-consistent attached eddy. W&M utilize an eddy that has a complex shape (Figure 1 in W&M), presumably configured using insight from DNS and/or PIV fields. Guided by the inferred kernels, we construct an attached-eddy template from a small number of Rankine vortex rods (a “square hairpin”) together with an inviscid image system. The objective is not an exact Navier–Stokes solution, but a kinematically consistent building block that (i) respects wall-impermeability, (ii) has well-posed planar influence integrals, and (iii) admits a transparent connection between morphology and the influence functions.
@@
    A spectral Influence kernel and interpretation of attached-eddy spectra. We introduce an explicit spectral Influence kernel $I_{\phi}(\kappa_{x},\eta)$ that maps a self-similar wall-parallel footprint to its one-dimensional streamwise energy spectrum as a function of relative height $\eta=y/h$. Combining $I_{\phi}$ with the scale-invariant population law $p(h)\propto h^{-3}$ yields a compact explanation for the emergence (and limitations) of the $k_{x}^{-1}$ range, the location of the premultiplied hump, and the origin of low-$k_{x}$ discrepancies under the independent-mark (Poisson) hypothesis..
All forward and inverse modeling tools used here are open sourced [^1] https://github.com/CaslabUM/AttachedEddy.
@@ 1  Statistics of attached eddies
Consider a collection of $n$ eddies (Figure 1) of length scale $h_{e,i}$ that are placed on the wall at locations $\mathbf{x}_{e,i}=(x_{e,i},z_{e,i})$. Since these are attached eddies, it is implicit that only the streamwise and spanwise components of $\mathbf{x}_{e,i}$ are variable. Define $\mathbf{h}_{e}\triangleq\{h_{e,1},h_{e,2},..,h_{e,n}\}$ and $\mathbf{X}\triangleq\{\mathbf{x}_{e,1},\mathbf{x}_{e,2},..,\mathbf{x}_{e,n}\}$. Assuming self-similarity and linearity, a quantity $q$ (e.g. a velocity component) evaluated at a location $\mathbf{x}$ can be determined using the superposition
@@
Homogenizing in the $(x,z)$ directions, and assuming $L$ is large enough that each eddy centered at the origin has a negligible induced contribution outside an area $4L^{2}$ (Refer Campbell’s theorem rice1944mathematical and the appendix of W&M), it can be shown that
@@
$=\beta\int_{h_{\min}}^{h_{\max}}p(h)\,h^{2}\left[\int_{-L/h}^{L/h}\int_{-L/h}^{L/h}u_{1}\!\left(\frac{\mathbf{x}}{h}\right)\,d\!\left(\frac{x}{h}\right)\,d\!\left(\frac{z}{h}\right)\right]dh+U_{ref}$
@@
The final piece we need is the probability distribution of the eddy sizes. Using insight from townsend1976structure and perry1982mechanism, W&M propose that $p(h)\propto 1/h^{3}$. In particular, for $h\in[h_{\min},h_{\max}]$,
@@
Here $h_{\max}\approx\delta^{+}$ is the outer length scale (boundary layer thickness in wall units) and $h_{\min}$ is set by the friction Reynolds number $Re_{\tau}$. For notational simplicity, all length and velocity scales should be assumed to be in wall units henceforth.
@@ 1.1  Statistical assumptions and derivation of the Reynolds-stress formula
    Independent marking: conditional on the locations, the marks $\{\theta_{k}\}$ are i.i.d. and independent of the point locations; in particular the size distribution is $p(h)$.
@@ 1.2  Convergence of planar influence integrals
Equation 4 defines the mean-flow eddy contribution (influence) function as a planar integral of the streamwise velocity induced by a single prototypical eddy. Writing the dimensionless coordinates $(\xi,\zeta,\eta)\triangleq(x/h,\;z/h,\;y/h)$ and $R\triangleq L/h$, we can make the dependence on the truncation radius explicit:
@@
whenever the limit exists. In particular, we do not assume compact support of the induced velocity field. Rather, the required condition for Eq. 4 to be well-defined is that the planar integral converges as $R\to\infty$, or equivalently that the tail contribution beyond $|(\xi,\zeta)|\gtrsim R$ becomes negligible for the chosen truncation.
$\int_{\mathbb{R}^{2}}|u_{1}(\xi,\zeta,\eta)|\,d\xi\,d\zeta<\infty.$ (13)
Reynolds-stress kernels are less restrictive.
Similarly, the second-moment influence functions are
For line-vortex Biot–Savart fields, a generic open filament can generate a $\rho^{-2}$ contribution to certain velocity components in the far field, which is precisely the borderline case for $I_{1}$ and would lead to a logarithmic growth of $I_{1}(\eta;R)$ with $R$. The specific square-hairpin eddy used herein is paired with an opposite-sign mirror image across $y=0$ (to enforce inviscid wall boundary conditions), and the wall-parallel segment cancels with its image. This construction removes leading-order contributions responsible for a $\rho^{-2}$ tail in the streamwise component and yields a faster far-field decay. In other words, the image construction is not a technical convenience: it is central to ensuring that Eq. 4 is well posed for a line-vortex eddy model.
@@ 2  The ideal eddy contribution function
$U(y)-U_{ref}$ $=\beta\int_{h_{\min}}^{h_{\max}}p(h)\,h^{2}\,I_{1}\!\left(\frac{y}{h}\right)\,dh,$ (16)
$R_{ij}(y)$ $=\beta\int_{h_{\min}}^{h_{\max}}p(h)\,h^{2}\,I_{ij}\!\left(\frac{y}{h}\right)\,dh.,$ (17)
where $U_{ref}$ is the outer velocity (e.g. center line or freestream velocity). For the classical attached-eddy size distribution $p(h)=C/h^{3}$, these become
$U(y)-U_{ref}=\beta\,C\int_{h_{\min}}^{h_{\max}}\frac{I_{1}(y/h)}{h}\,dh,\qquad R_{ij}(y)=\beta\,C\int_{h_{\min}}^{h_{\max}}\frac{I_{ij}(y/h)}{h}\,dh,$ (18)
@@
$U(y)-U_{ref}=\beta\,C\int_{y/h_{\max}}^{y/h_{\min}}\frac{I_{1}(\eta)}{\eta}\,d\eta,\qquad R_{ij}(y)=\beta\,C\int_{y/h_{\max}}^{y/h_{\min}}\frac{I_{ij}(\eta)}{\eta}\,d\eta.$ (19)
@@ 2.1  Discrete formulation and identifiability
$\mathbf{U}(\tilde{\mathbf{y}})-U_{ref}\mathbf{1}=\beta\,\mathbf{K}(\mathbf{c})\mathbf{b},$ (20)
@@
$U(y_{r};\mathbf{d})$ $=\beta\sum_{l=1}^{m}p(h_{l})h_{l}^{2}\left[\sum_{q}\sum_{p}u_{1}\left(\frac{{x_{p},z_{q},y_{r}}}{h_{l}};\mathbf{d}\right)\left(\frac{\Delta x}{h_{l}}\right)\left(\frac{\Delta z}{h_{l}}\right)\right]\Delta h+U_{ref}$
$R_{ij}(y_{r};\mathbf{d})$ $=\beta\sum_{l=1}^{m}p(h_{l})h_{l}^{2}\left[\sum_{q}\sum_{p}u_{i}\left(\frac{{x_{p},z_{q},y_{r}}}{h_{l}};\mathbf{d}\right)u_{j}\left(\frac{{x_{p},z_{q},y_{r}}}{h_{l}};\mathbf{d}\right)\left(\frac{\Delta x}{h_{l}}\right)\left(\frac{\Delta z}{h_{l}}\right)\right]\Delta h.$
Explicitly $K_{ij}=I_{1}(z_{i}/h_{j})$ and $b_{i}=\beta p(h_{i})h_{i}^{2}\Delta h.$ We would now like to extract these coefficients. Towards this end, we define the eddy contribution function at selected locations, and interpolate for the values in other locations in a piecewise linear fashion. In other words
@@
$\triangleq\begin{bmatrix}c_{1}&c_{m+1}&c_{m+2}&..&c_{2m-2}&c_{2m-1}\\ c_{2}&c_{1}&I_{1}(h_{2}/h_{3})&..&I_{1}(h_{2}/h_{m-1})&I_{1}(h_{2}/h_{m})\\ ..&..&..&..&..&..\\ c_{m}&I_{1}(h_{m}/h_{2})&I_{1}(h_{m}/h_{3})&..&I_{1}(h_{m}/h_{m-1})&c_{1}\\ \end{bmatrix}$.
The unknown $I_{1}(\cdot)$ values are then interpolated  [^2] For instance, if $h_{\min}=186$ and $h_{\max}=5186$ and $m=11$, then $I_{1}(h_{2}/h_{3})=0.42106c_{1}+0.57894c_{12}$ from the nodal locations $\mathbf{c}$. Note that there are more unknowns ($2m$) than equations ($m$), and so while it is possible to determine the unknowns that lead to perfect match of a reference (DNS or experiment) velocity profiles, a choice has to be made on the problem formulation. We choose to solve the following minimum norm least squares problem:
$\mathbf{c}_{\textrm{opt}}=\min\mathbf{c}^{T}\mathbf{c}\ \ \textrm{such that}\ \ \mathbf{U}_{ref}(\tilde{\mathbf{y}})-U_{ref}=\mathbf{K}\mathbf{b}.$
Figure 2 shows the optimal eddy contribution functions inferred from the channel flow data (lee2015direct) at friction Reynolds number $Re_{\tau}\approx 5200$. Figure 3 confirms that the first and second velocity moments are perfectly reproduced across the entire channel.
@@
Figure 3: Reference (symbols) vs optimal attached eddy statistics for $Re_{\tau}\approx 5200$

### v5

@@ Abstract
Townsend’s attached eddy hypothesis models the logarithmic region of high Reynolds number wall turbulence as a random superposition of wall-attached, geometrically self-similar eddies whose sizes obey a scale-invariant population law. Building on the statistical framework of Woodcock & Marusic (2015), the present work (i) poses an inverse problem to infer the ideal single-eddy contribution (influence) functions for the mean velocity and Reynolds stresses from DNS moments, (ii) uses these inferred kernels to guide a minimal Biot–Savart-consistent hairpin-type eddy built from Rankine vortex rods together with an inviscid image system, and (iii) introduces and infers a spectral Influence kernel that maps a self-similar eddy footprint to its one-dimensional energy spectrum. The Influence-kernel viewpoint yields a transparent explanation for the emergence (and limitations) of the linear part of the energy spectrum, provides a clear scale-by-scale decomposition and helps rationalize why simple eddy templates can reproduce a broad set of log-layer statistics once the mean-flow anchoring is fixed. Exact closed-form expressions for the mean influence function and the Fourier-space streamwise velocity of a general straight-segment hairpin family with image are derived, revealing a clean mean-variance duality: the horizontal head determines the entire mean kernel $I_{1}$ while the inclined legs dominate the spectral energy $I_{\phi}$. This structural insight explains why the rectangular hairpin occupies a singular corner of the eddy design space and why replacing it without degrading either mean or spectral predictions is difficult. The model is further extended by allowing the eddy population density to vary with scale, yielding near-perfect predictions of mean velocity and streamwise variance across $Re_{\tau}=6000$2000020000.
@@
Characterizing and predicting the turbulent flow in the vicinity of a wall continues to not only be a problem of scientific interest, but also one of great practical relevance. Significant knowledge and insight has been gained from the development of experimental diagnostics in the 1960s to the emergence of direct numerical simulations in the 1980s, sophisticated measurements in the 1990s, an emphasis on coherent structures in the 2000s and operator-based dynamical models over the past decade. Against this backdrop of increasingly sophisticated experimentation, computations and analyses, the attached eddy hypothesis - originally proposed more than 50 years ago by  Townsend (1976) - continues to serve as a simple, yet effective theory to qualitatively and quantitatively describe structural and statistical aspects of turbulent boundary layers, particularly focusing on the logarithmic region.
The attached eddy model (AEM) was given a firm mathematical footing by  Perry & Chong (1982), and further developed by Marusic and co-workers over the past 25 years  (Perry & Marušić, 1995; Marusic et al., 2013; Woodcock & Marusic, 2015; de Silva et al., 2016b). Among other useful features, the AEM is able to explain scaling behaviors of velocity moments, provide an explanation for uniform momentum zones (de Silva et al., 2016a), and serve as a predictive model for the von Kármán constant as a function of Reynolds number. A review of the theory and developments of the AEM of wall turbulence can be found in  Marusic & Monty (2019). Over the past few years, theoretical and numerical analyses (e.g.  McKeon (2019); Lozano-Durán & Bae (2019)) have added further credibility to this theory. Despite the success of the AEM, it is pertinent to remember that it is fundamentally a statistical theory, and other hypotheses (e.g.  Davidson et al. (2006); Davidson & Krogstad (2009)) can also be used to explain the statistical characteristics of the log layer.
@@
A key advance was made by Woodcock & Marusic (2015) (henceforth W&M) who established a statistical foundation for AEM. They provide a complete derivation for all the velocity moments and demonstrate logarithmic scaling relationships therein. They were also able to provide expressions for the skewness and flatness of the wall-normal and spanwise fluctuations as a function of the Reynolds number. While variants of the AEM continue to be developed in the literature (e.g. Hwang & Eckhardt (2020)), we consider W&M as the starting point of our exploration and exclusively consider zero pressure gradient boundary layers. The main contributions and guiding questions in this work are:
@@
    Inverse identification of attached-eddy influence functions. Townsend introduced the notion of an “eddy intensity” (or contribution) function: a wall-parallel average of the single-eddy induced velocity and its products. While Townsend gave remarkably insightful descriptions of the nature of this function (Figure 5.7, page 155 of Townsend (1976)), his and most other treatments are qualitative. Here we pose a concrete inverse problem that infers the ideal influence functions $I_{1}(y/h)$ and $I_{ij}(y/h)$ implied by reference one-point moments (DNS/experiment). The inferred kernels expose the wall-normal support, sign structure, and plateau/decay behavior required for log-layer mean and Reynolds stresses, and motivate a simplified kernel model that yields a direct relation between the von Kármán constant $\kappa$ and the population density $\beta$.
@@
    A minimal Biot–Savart-consistent attached eddy. W&M utilize an eddy that has a complex shape (Figure 1 in W&M), presumably configured using insight from DNS and/or PIV fields. Guided by the inferred kernels, we construct an attached-eddy template from a small number of Rankine vortex rods (a “rectangular hairpin”) together with an inviscid image system. The objective is not an exact Navier–Stokes solution, but a kinematically consistent building block that (i) respects wall-impermeability, (ii) has well-posed planar influence integrals, and (iii) admits a transparent connection between morphology and the influence functions.
@@
    A spectral Influence kernel and interpretation of attached-eddy spectra. We introduce an explicit spectral Influence kernel $I_{\phi}(\kappa_{x},\eta)$ that maps a self-similar wall-parallel footprint to its one-dimensional streamwise energy spectrum as a function of relative height $\eta=y/h$. Combining $I_{\phi}$ with the scale-invariant population law $p(h)\propto h^{-3}$ yields a compact explanation for the emergence (and limitations) of the $k_{x}^{-1}$ range, the location of the premultiplied hump, and the origin of low-$k_{x}$ discrepancies under the independent-mark (Poisson) hypothesis. The ideal influence kernel is extracted by defining a Fredholm-type inverse problem.
  5. 5.
    Why the rectangular hairpin is unusually predictive. Exact closed-form expressions for the mean kernel and Fourier-space velocity of a general three-segment hairpin family reveal a clean mean-variance duality: the horizontal head determines the entire mean kernel $I_{1}$, while the inclined legs dominate the spectral energy $I_{\phi}$. Only the rectangular hairpin ($b_{1}=b_{0}$) produces an exact plateau in $I_{1}$ and hence a purely logarithmic mean. This structural insight explains why it occupies a singular corner of the eddy design space and why replacing it without degrading either mean or spectral predictions is difficult.
Forward and inverse modeling tools used here are open sourced [^1] https://github.com/CaslabUM/AttachedEddy.
@@ 1  Statistics of attached eddies
Consider a collection of $n$ eddies (Figure 1) of length scale $h_{e,i}$ that are placed on the wall at locations $\mathbf{x}_{e,i}=(x_{e,i},z_{e,i})$. Since these are attached eddies, it is implicit that only the streamwise and spanwise components of $\mathbf{x}_{e,i}$ are variable. Define $\mathbf{h}_{e}\triangleq\{h_{e,1},h_{e,2},..,h_{e,n}\}$ and $\mathbf{X}_{e}\triangleq\{\mathbf{x}_{e,1},\mathbf{x}_{e,2},..,\mathbf{x}_{e,n}\}$. Assuming self-similarity and linearity, a quantity $q$ (e.g. a velocity component) evaluated at a location $\mathbf{x}$ can be determined using the superposition
@@
Homogenizing in the $(x,z)$ directions, and assuming $L$ is large enough that each eddy centered at the origin has a negligible induced contribution outside an area $4L^{2}$ (cf. Campbell’s theorem Rice (1944) and the appendix of W&M), it can be shown that
@@
The final piece we need is the probability distribution of the eddy sizes. Using insight from Townsend (1976) and Perry & Chong (1982), W&M propose that $p(h)\propto 1/h^{3}$. In particular, for $h\in[h_{\min},h_{\max}]$,
@@
Here $h_{\max}\approx\delta^{+}$ is the outer length scale (boundary layer thickness in wall units) and $h_{\min}$ is set by the friction Reynolds number in the form $2.6\sqrt{Re_{\tau}}$.
@@ 1.1  Statistical assumptions and derivation of the Reynolds-stress formula
    Independent marking: conditional on the locations, the marks $\{\theta_{k}\}$ (collectively, all eddy properties beyond the wall-parallel center location) are i.i.d. and independent of the point locations; in particular the size distribution is $p(h)$.
@@ 1.2  Convergence of planar influence integrals
Equation (4) defines the mean-flow eddy contribution (influence) function as a planar integral of the streamwise velocity induced by a single prototypical eddy. Writing the dimensionless coordinates $(\xi,\zeta,\eta)\triangleq(x/h,\;z/h,\;y/h)$ and $R\triangleq L/h$, we can make the dependence on the truncation radius explicit:
@@
whenever the limit exists. In particular, we do not assume compact support of the induced velocity field. Rather, the required condition for Eq. (4) to be well-defined is that the planar integral converges as $R\to\infty$, or equivalently that the tail contribution beyond $|(\xi,\zeta)|\gtrsim R$ becomes negligible for the chosen truncation. A detailed analysis of convergence is presented in Appendix A.
@@ 2  The ideal eddy contribution function
$U_{\rm data}(y)-U_{ref,\rm{data}}$ $=\beta\int_{h_{\min}}^{h_{\max}}p(h)\,h^{2}\,I_{1}\!\left(\frac{y}{h}\right)\,dh,$ (13)
$R_{ij,\rm{data}}(y)$ $=\beta\int_{h_{\min}}^{h_{\max}}p(h)\,h^{2}\,I_{ij}\!\left(\frac{y}{h}\right)\,dh.$ (14)
where $U_{ref,\rm{data}}$ is the outer velocity (e.g. center line or freestream velocity). For the classical attached-eddy size distribution $p(h)=C/h^{3}$, these become
$U_{\rm data}(y)-U_{ref,\rm{data}}=\beta\,C\int_{h_{\min}}^{h_{\max}}\frac{I_{1}(y/h)}{h}\,dh,\qquad R_{ij,\rm{data}}(y)=\beta\,C\int_{h_{\min}}^{h_{\max}}\frac{I_{ij}(y/h)}{h}\,dh,$ (15)
@@
$U_{\rm data}(y)-U_{ref,\rm{data}}=\beta\,C\int_{y/h_{\max}}^{y/h_{\min}}\frac{I_{1}(\eta)}{\eta}\,d\eta,\qquad R_{ij,\rm{data}}(y)=\beta\,C\int_{y/h_{\max}}^{y/h_{\min}}\frac{I_{ij}(\eta)}{\eta}\,d\eta.$ (16)
@@ 2.1  Discrete formulation and identifiability
$\mathbf{U}_{\rm data}(\tilde{\mathbf{y}})-U_{ref,\rm{data}}\mathbf{1}=\mathbf{K}(\mathbf{c})\mathbf{b},$ (17)
@@
$U_{\rm data}(y_{r};\mathbf{d})$ $=\beta\sum_{l=1}^{m}p(h_{l})h_{l}^{2}\left[\sum_{q}\sum_{p}u_{1}\left(\frac{{x_{p},z_{q},y_{r}}}{h_{l}};\mathbf{d}\right)\left(\frac{\Delta x}{h_{l}}\right)\left(\frac{\Delta z}{h_{l}}\right)\right]\Delta h+U_{ref,\rm{data}}$
$R_{ij,\rm{data}}(y_{r};\mathbf{d})$ $=\beta\sum_{l=1}^{m}p(h_{l})h_{l}^{2}\left[\sum_{q}\sum_{p}u_{i}\left(\frac{{x_{p},z_{q},y_{r}}}{h_{l}};\mathbf{d}\right)u_{j}\left(\frac{{x_{p},z_{q},y_{r}}}{h_{l}};\mathbf{d}\right)\left(\frac{\Delta x}{h_{l}}\right)\left(\frac{\Delta z}{h_{l}}\right)\right]\Delta h.$
Explicitly $K_{ij}=I_{1}({y}_{i}/h_{j})$ and $b_{i}=\beta p(h_{i})h_{i}^{2}\Delta h.$ We would now like to extract these coefficients. Towards this end, we define the eddy contribution function at selected locations, and interpolate for the values in other locations in a piecewise linear fashion. In other words
@@
$\triangleq\begin{bmatrix}c_{1}&c_{m+1}&c_{m+2}&..&c_{2m-2}&c_{2m-1}\\ c_{2}&c_{1}&c_{m+1}&..&c_{2m-3}&c_{2m-2}\\ ..&..&..&..&..&..\\ c_{m}&c_{m-1}&c_{m-2}&..&c_{2}&c_{1}\\ \end{bmatrix}$.
The unknown $I_{1}(\cdot)$ values are then interpolated  [^2] For instance, if $h_{\min}=186$ and $h_{\max}=5186$ and $m=11$, then $I_{1}(h_{2}/h_{3})=0.42106c_{1}+0.57894c_{12}$ from the nodal locations $\mathbf{c}$. Note that there are more unknowns ($2m$) than equations ($m$), and so while it is possible to determine the unknowns that lead to perfect match of a reference (DNS or experiment) velocity profiles, a choice has to be made on the problem formulation. We choose to solve the following minimum-norm exact interpolation problem:
$\mathbf{c}_{\textrm{opt}}=\min\mathbf{c}^{T}\mathbf{c}\ \ \textrm{such that}\ \ \mathbf{U}_{\rm{data}}(\tilde{\mathbf{y}})-U_{ref,\rm{data}}=\mathbf{K}\mathbf{b}.$
Figure 2 shows the optimal eddy contribution functions inferred from the channel flow data (Lee & Moser, 2015) at friction Reynolds number $Re_{\tau}\approx 5200$. Figure 3 confirms that the first and second velocity moments are perfectly reproduced across the entire channel. It is emphasized that the perfect reproduction in Figure 3 is expected by construction: the system is underdetermined ($2m$ unknowns, $m$ equations), and the minimum-norm solution exactly satisfies the constraints. The sharp features near $\eta\approx 1$ visible in Figure 2 are sensitive to the discretization and the minimum-norm selection criterion; by contrast, the qualitative structure (a near-plateau for $\eta\lesssim 1$ followed by rapid decay) is robust across different discretizations, regularization strategies, and data sources. It is this robust qualitative structure, rather than the fine details, that motivates the simplified model in Section 2.2. The inversion uses moments across the full channel half-height ($y^{+}$ from $h_{\min}$ to $h_{\max}$), so the inferred $I_{1}(\eta)$ and $I_{ij}(\eta)$ are constrained by the entire wall-normal profile, not only the log layer.
@@
Figure 3: Reference (symbols) vs optimal attached eddy statistics for $Re_{\tau}\approx 5200$. The moments are reconstructed from the inferred $I_{1}$ and $I_{ij}$ via Eqs. (3) and (11). The lowest $y^{+}$ shown corresponds approximately to $h_{\min}$, below which the AEM has no eddies and the model is not expected to apply.